# Code Readability in the Age of Large Language Models: An Industrial Case Study from Atlassian

### v1

@@ 4.2. Results
Figure 2 shows the results of this research question, covering the case study of HULA on our internal dataset. Table 5 shows the statistical significance and the effect size of the result.
Our case study on the HULA framework reveals that the generated code is comparable to the human-written code, with only minor differences. Figure 2 shows that HULA, with a GPT-4 based model, generates code that is highly similar to humans with only a slightly longer code but with no statistically significant difference in Cyclomatic Complexity and comment ratio. On Halstead’s metrics, our LLM-generated code has marginally higher values on difficulty, vocabulary, volume, and time required. These results indicate that the HULA framework can produce code that closely resembles human-written code in terms of code readability, indicating the promising performance of LLM in production.
Table 5. (RQ2) The mean significance from Mann-Whitney U Test (Mann and Whitney, 1947) and Effect Size measured by the Cliff’s Delta (Cliff, 1993) based on the comparison of HULA and Human code.
Metric P-Value Effect Size Line of Code <0.001 0.230 (small) Comment Ratio 0.005 0.124 Cyclomatic Complexity 0.622 0.015 Maintainability Index <0.001 -0.225 (small) Halstead Difficulty <0.001 0.201 (small) Halstead Vocabulary <0.001 0.202 (small) Halstead Volume <0.001 0.208 (small) Halstead Time <0.001 0.205 (small)
[…]
[Code Readability Measurement]The Scores of Code Readability Measurement between Human and HULA Code on the Internal Dataset.

### v2

@@ 4.2. Results
Figure 5 shows the results of this research question, covering the case study of HULA on our internal dataset. Table 8 shows the statistical significance and the effect size of the result. Lastly, Figure 6 shows the detailed analysis of code readability measurement by programming languages.
The generated code is comparable to the human-written code with a negligible to small statistical difference. Figure 5 shows that HULA, with a GPT-4 based model, generates code that is highly similar to humans with only a slightly longer code but with no statistically significant difference in Cyclomatic Complexity and comment ratio. On Halstead’s metrics, our LLM-generated code has marginally higher values on difficulty, vocabulary, volume, and time required. These results indicate that the HULA framework can produce code that closely resembles human-written code in terms of code readability, indicating the promising performance of LLM in production.
The HULA-generated code is comparable to the human-written code specifically in Java, Kotlin, Go and Scala programming languages. Figure 6 shows the static analysis results by programming languages of code generated by Hula and written by humans. The case study on the Hula framework reveals two different patterns across the six studied programming languages. First, in Java, Kotlin, Go and Scala programming languages, HULA-generated code shows comparable results across all code readability measures to that written by humans. Second, in TypeScript and Python programming languages, the Hula-generated code tends to have longer lines of code and slightly lower maintenance scores than the human-written code by small effect sizes.
Table 8. (RQ2) The mean significance from Mann-Whitney U Test (mann1947test) and Effect Size measured by the Cliff’s Delta (cliff1993dominance) based on the comparison of HULA and Human code.
Metric P-Value Effect Size
Line of Code <0.001 0.230 (small)
Comment Ratio 0.005 0.124 (negligible)
Cyclomatic Complexity 0.622 0.015 (negligible)
Maintainability Index <0.001 -0.225 (small)
Halstead Difficulty <0.001 0.201 (small)
Halstead Vocabulary <0.001 0.202 (small)
Halstead Volume <0.001 0.208 (small)
Halstead Time <0.001 0.205 (small)
@@ 5. Threat to Validity
Additionally, the empirical study was conducted on code generation tasks using GPT-4 as a based model and the Atlassian internal dataset. Efforts were made to ensure diversity in programming languages and task contexts. However, it is acknowledged that the findings may vary when applied to different datasets. Therefore, future research is encouraged to investigate code readability across a broader range of contexts.
Lastly, while the HULA framework typically operates with a human-in-the-loop—where software engineers provide feedback to guide the outputs of the LLM-based agent, in this study, all human feedback and intervention were intentionally excluded. The effort was made to avoid introducing additional information to the LLM-based agents and to ensure a fair comparison between code written by humans and code generated by LLM-based agents. Future studies could compare results from both autonomous and human-in-the-loop settings to better understand their respective trade-offs.
@@
Last but not least, although we may not be able to disclose the source code of the evaluation dataset due to industrial confidentiality reasons, we are confident that the evaluation offers valuable insight into the practical use of LLMs in real-world development workflows and reflects industrial standards for production-level code, paving the way forward the future of AI-powered software development tools.